# Quantitative classification of potential Navier-Stokes singularities beyond the blow-up time

### v1

@@ Abstract.
Navier-Stokes equations, axisymmetric solutions, local quantitative estimates, blow-up profile, localized energy norm
@@ 1.1. Main results
(1.10) $\int\limits_{M^{c_{0}(p)}T^{\frac{1}{2}}<|y|<M^{c_{1}(p)}T^{\frac{1}{2}}}|v(y,T)|^{2}dy\geq T^{\frac{1}{2}}e^{-M^{c_{2}(p)}}\qquad\forall T\in[T^{*},T^{*}(1+\lambda_{0})].$
@@
(1.12) $\int\limits_{M^{c_{3}(p)}<|y|<e^{M^{c_{4}(p)}}}|v(y,T)|^{2}dy\geq e^{-e^{M^{c_{5}(p)}}}\qquad\forall T\in[M^{-5},M^{5}].$
@@
(1.13) $\int\limits_{M^{c_{0}(p)}T^{\frac{1}{2}}<|y|<M^{c_{1}(p)}T^{\frac{1}{2}}}|v(y,T)|^{2}dy\geq T^{\frac{1}{2}}e^{-M^{c_{2}(p)}}\qquad\forall T\in[M^{-5},T^{*}(1+\lambda_{0})].$
@@ 1.3.1. Optimality
$\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|}|e^{T\Delta}v_{0}(x)|^{2}dx\leq T^{\frac{1}{2}}e^{-M^{1400}}.$
@@
$\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|^{2}dx\geq T^{\frac{1}{2}}e^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[T^{*},T^{*}(1+\lambda_{0})],\end{split}$
@@ 1.3.2. Vertical decay on the initial data
1.3.3. Initial data that is not approximately axisymmetric
respectively.
@@ 5.1. Proof of Theorem 1
(5.7) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|^{2}dx\geq T^{\frac{1}{2}}e^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[T^{*},T^{*}(1+\lambda_{0})].\end{split}$
@@ 5.2.1. Case 1: $T\in[M^{-5},T^{*}(1+\lambda_{0})]$
(5.9) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|^{2}dx\geq T^{\frac{1}{2}}e^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[M^{-12}T^{*},T^{*}].\end{split}$
@@
(5.10) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|^{2}dx\geq T^{\frac{1}{2}}e^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[M^{-5},T^{*}(1+\lambda_{0})].\end{split}$
@@ Acknowledgments
The author would like to thank Alexey Cheskidov and László Székelyhidi for interesting discussions before the genesis of this paper. The author is supported by an EPSRC New Investigator Award.

### v2

@@ Abstract.
Navier-Stokes equations, axisymmetric solutions, local quantitative estimates, blow-up profile, localized norm
@@ 1.1. Main results
(1.10) $\int\limits_{M^{c_{0}(p)}T^{\frac{1}{2}}<|y|<M^{c_{1}(p)}T^{\frac{1}{2}}}|v(y,T)|dy\geq Te^{-M^{c_{2}(p)}}\qquad\forall T\in[T^{*},T^{*}(1+\lambda_{0})].$
@@
(1.12) $\int\limits_{M^{c_{3}(p)}<|y|<e^{M^{c_{4}(p)}}}|v(y,T)|dy\geq e^{-e^{M^{c_{5}(p)}}}\qquad\forall T\in[M^{-5},M^{5}].$
@@
(1.13) $\int\limits_{M^{c_{0}(p)}T^{\frac{1}{2}}<|y|<M^{c_{1}(p)}T^{\frac{1}{2}}}|v(y,T)|dy\geq Te^{-M^{c_{2}(p)}}\qquad\forall T\in[M^{-5},T^{*}(1+\lambda_{0})].$
@@ 1.3. Remarks
1.3.1. Initial data that is not approximately axisymmetric
respectively.
@@ 1.3.2. Optimality
$\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|}|e^{T\Delta}v_{0}(x)|dx\leq Te^{-M^{1400}}.$
@@
$\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|dx\geq Te^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[T^{*},T^{*}(1+\lambda_{0})],\end{split}$
@@ 5.1. Proof of Theorem 1
(5.7) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|dx\geq Te^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[T^{*},T^{*}(1+\lambda_{0})].\end{split}$
@@ 5.2.1. Case 1: $T\in[M^{-5},T^{*}(1+\lambda_{0})]$
(5.9) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|dx\geq Te^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[M^{-12}T^{*},T^{*}].\end{split}$
@@
(5.10) $\begin{split}\int\limits_{M^{708+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}\leq|x|\leq M^{1102+\frac{30}{1-\frac{3}{p}}}T^{\frac{1}{2}}}&|v(x,T)|dx\geq Te^{-M^{1813+\frac{60}{1-\frac{3}{p}}}}\\ &\forall T\in[M^{-5},T^{*}(1+\lambda_{0})].\end{split}$
@@ Acknowledgments
The author would like to thank Alexey Cheskidov and László Székelyhidi for interesting discussions before the genesis of this paper. The author is supported by the EPSRC New Investigator Award UKRI096 ‘Dynamics and regularity criteria for nonlinear incompressible partial differential equations’.